OpenStudy (anonymous):

What is the solution of |2x-1|=5

OpenStudy (anonymous):

x=-2 and x=3

OpenStudy (anonymous):

|2x-1|=5 2x-1=5 or 2x-1=-5 2x-1=5 2x=6 x=3 2x-1=-5 2x=-4 x=-2

OpenStudy (anonymous):

thank you.